Mathematics 242 – Principles of Analysis Definitions Quiz 6 – March 25, 2011

Your Name:

A) (10) What does it mean for a sequence {x_{n}} to converge to a ∈ R?

B) (10) Give the ε, δ-version of the definition of the statement that f is continuous at x = c.